OpenStudy (anonymous):

Help please! P(x)=150(1.07)^x where x is the number of years since 1995. in which year will the population reach 300

OpenStudy (anonymous):

a)2005 b)2028 c)2016 d)2026

OpenStudy (anonymous):

sorry b is incorrect its 2038

OpenStudy (campbell_st):

well let P(x) =300 so \[300 = 150(1.07)^x\] divide both sides of the equation by 150. take the log of both sides this should make it easier for find x, you'll just need to apply the log law for powers \[\log(x^a) = alog(x)\]

OpenStudy (anonymous):

i got 2005?

OpenStudy (campbell_st):

seems a sensible answer